i. If the price of rice per Kg increases from Rs 200 to Rs 300, the quantity demand reduced from 10 Kg units to 8 Kg. Keeping the answer in view if price change by 7%, what happen to Quantity demand of Rice. Give suitable example.
1
Expert's answer
2020-12-17T07:42:15-0500

"\u2206in price = 300-200 = Rs100""\u2206 in quantity= 8-10= -2kg" "Percentage change in price=100\/200\u00d7100=50%""7\u00d7100\/50=Rs14"

An increase in price results to a decrease in quantity of good demanded. After an increase in price by Rs 100 the quantity of good demanded decreased by-2kg

An increase of goods demanded by 7% results to a decrease in price by Rs14